OpenStudy (anonymous):

graph the following quadratic equation: y+12+x^2=4x^2-7x

OpenStudy (mertsj):

y=3x^2-7x-12

OpenStudy (anonymous):

ok i have that but i dont know how to graph it '

OpenStudy (mertsj):

x=-b/2a=7/6. So the x coordinate of the vertex is 7/6. Now plug 7/6 in to find the y coordinate of the vertex

OpenStudy (mertsj):

So the vertex is (7/6,-14 11/12)

OpenStudy (mertsj):

Vertex is (7/6,-16 1/12)

OpenStudy (mertsj):

Now choose an x value to the right of 7/6 like 2 and plug it in to find y.

OpenStudy (mertsj):

You will get (2,-14)

OpenStudy (mertsj):

Use symmetry to get the corresponding point to the left of the vertex.

OpenStudy (mertsj):

Now choose another x value perhaps 3 and plug it in. You will get -6. So graph (3,-6). Graph the corresponding point to the left of the vertex. That should be enough to draw the curve.

OpenStudy (anonymous):

i am so lost'

OpenStudy (anonymous):

Two versions of the plotted quadratic expression is attached.

OpenStudy (anonymous):

ok thank you

OpenStudy (mertsj):

So money, do you still need help?

OpenStudy (mertsj):

I see you said you are so lost.

OpenStudy (anonymous):

no i got it

OpenStudy (mertsj):

ok